OpenStudy (driftracer305):

what is the standard form equation of the ellipse with vertices at (0 6) and (0 -6) and co-vertices at (4 0) and (-4 0)..?

OpenStudy (dumbcow):

 standard form: \[\frac{y^2}{a^2} + \frac{x^2}{b^2} = 1\] Note: "a" is under the y term because it is length of vertical axis

OpenStudy (driftracer305):

sweet @dumbcow sir..... but just concerns me as that's not in my options

OpenStudy (dumbcow):

well now thats no good a^2 = 36 b^2 = 16 right

OpenStudy (driftracer305):

yes

OpenStudy (driftracer305):

i know it has to be B.... right @dumbcow ?

OpenStudy (dumbcow):

yes i just switched the order ... yeah standard form wants the "x" to come first

OpenStudy (driftracer305):

k thank u sir......